### v1

# From Radiation Dose to Cellular Dynamics: A Discrete Model for Simulating Cancer Therapy

Abstract
Radiation therapy is one of the most common cancer treatments, and dose optimization and targeting of radiation are crucial since both cancerous and healthy cells are affected. Different mathematical and computational approaches have been developed for this task. The most common mathematical approach, dating back to the late 1970’s, is the linear-quadratic (LQ) model for the survival probability given the radiation dose. Most simulation models consider tissue as a continuum rather than consisting of discrete cells. While reasonable for large scale models, e.g., for human organs, any cellular scale effects become, by necessity, neglected. They do, however, influence growth, morphology, and metastasis of tumors. Here, we propose a method for modeling the effect of radiation on cells based on the mechanobiological CellSim3D simulation model for growth, division, and proliferation of cells. To model the effect of a radiation beam, we incorporate a Monte Carlo procedure into CellSim3D with the LQ model by introducing a survival probability at each beam delivery. Effective removal of dead cells by phagocytosis was also implemented. Systems with two types of cells were simulated: stiff slowly proliferating healthy cells and soft rapidly proliferating cancer cells. For model verification, the results were compared to prostate cancer (PC-3 cell line) data for different doses and we found good agreement. In addition, we simulated proliferating systems and analyzed the probability density of the contact forces. We determined the state of the system with respect to the jamming transition and found very good agreement with experiments.
I Introduction
According to the US Centers for Disease Control and Prevention (CDC), cancer was the second most common cause of death responsible for 21.9% and 20.7% of deaths of males and females, respectively, in 2017 [1, 2]. It has been estimated that as many as 70% of cancer patients receive radiotherapy either as the main treatment or in combination with others [3]. Radiation treatment is particularly effective in breast and prostate cancers, and according to Bryant et al. [4], 40% of breast and 23% of prostate cancer survivors received radiation treatment.
Radiotherapy can be described as a treatment with the aim to dysfunctionalize cancer cell DNA by damaging it by irradiation of targeted high-energy beam of, e.g., x-rays, protons or electrons. Radiation is not selective and also damages healthy cells and their DNA, and thus may, in addition to immediate side effects, have side effects that take long times to be manifest. Due to the above reasons, optimization of spatial accuracy, dose and treatment frequency are critical for success. Reviews of the perspectives and history of radiotherapy are provided, for example, by Bernier et al. [5], Connell and Hellman [6], and Streffer and Herrmann [7]. Different approaches to modeling radiotherapy are provided, e.g., Jones and Dale [8].
There are different approaches to treatment planning algorithms. They are typically divided into correction-based and model-based, the latter being generally more accurate [9, 10]. Difficulties include, for example, inhomogeneous tissue, fractonation of dose, over-expression of self-repair mechanisms in cancer cells and cell migration. One further issue is that most models treat the tissue as a continuum rather than consisting of cells of discrete sizes and varying mechanical properties. It has recently been shown that the local mechanical stress landscape has a strong influence on cell migration, and that targeting mechanical properties may provide a new pathway to cancer treatment [11].
Some of the current modeling approaches include agent-based ones [12] which may offer a new and adaptable way for modeling; although not specific to radiotherapy, many of the current models for cancer are discussed in the editorial by Enderling and Rejniak [13] and other articles in that same special issue. Other reviews discussing radiotherapy related matters are provided, e.g., by Marcu and Harriss-Phillips [14], and D’Andrea et al. [15] As the reviews show, discreteness of cells is very rarely accounted for in radiotherapy modeling. One of the models in that direction is the one by Bobadilla et al. [16] where cells are differentiated according to whether they have received radiation therapy or not. Conceptually, we combine different levels of modeling using a mechanobiological cell model together with the continuum LQ model. This is in the same sprit as the work of Liu et al. [17], who combined radiation transport using the Geant4 framework, and a cell simulation platform called CompuCell3D.
The aim of this paper is to develop a radiotherapy model based on discrete cells with heterogeneous properties, each capable of responding to changes in its local environment. To this end, we used the CellSim3D model [18] as the basis, and integrated features of an external radiation beam by combining it with a Monte Carlo process simulation and the LQ approach. Using CellSim3D as the basis also provides an easily expandable tool beyond LQ. It is important to note that the current implementation assumes a spatially uniform or photon-based radiation field, which aligns with standard clinical photon therapies. Extending this framework to account for more complex radiation types, such as high-energy protons or heavier ions, remains an open and technically challenging problem. Preliminary studies in this direction exist [19, 20], particularly for high-energy protons where the biological effects tend to resemble those of photons under certain conditions. Nonetheless, accurately modeling the stochastic and spatially heterogeneous energy deposition characteristic of particle therapies would require significant advancements in both physical modeling and biological response integration, and represents a promising avenue for future work.
Simulation | Total # of steps ( | Write frequency () | Box side | Beam | Dose [Gy] | [Gy-1] | [Gy-2] |
Prostate | 7.25 | 2.5 | 330 | Present | 3 | 0.064 | 0.0167 |
6 | |||||||
9 | |||||||
15 | |||||||
20 | |||||||
3 | 0.241 | 0.067 | |||||
6 | |||||||
9 | |||||||
15 | |||||||
20 | |||||||
3 | 0.487 | 0.055 | |||||
6 | |||||||
9 | |||||||
15 | |||||||
20 | |||||||
Force distribution | 5.5 | 2 | 47 | Absent | 0 | 1 | 1 |
Present | 0.2333 | ||||||
0.4712 | |||||||
0.7792 | |||||||
500 | Absent | 0 | |||||
Present | 0.2333 | ||||||
0.4712 | |||||||
0.7792 |
II Methods
II.1 CellSim3D and inclusion of radiation beam
The CellSim3D [18] model and software for cell division was used as the starting point. CellSim3D is an open source software for mechanobiological simulations of cells populations [22]. We omit the model details here as they have been described in detail in several prior publications and since the code and model are freely available [18, 23, 24]; in particular, Supplemental Material provided in Ref. [24] has a very detailed description of the forces and interactions in the model.
In brief, CellSim3D is GPU accelerated and able to routinely handle 100,000’s of cells in 3-dimensions on a regular desktop computer equipped with standard NVidia graphics cards (a wide range has been tested as reported in Ref. [18]). The detailed derivation of the model and a description of the software can be found in Ref. [18] while the mapping between physical and model parameters is provided in Ref. [25]. The parameters used in this work are listed in Table 1 of Ref. [18] and will not be rewritten here. The model includes cell-cell adhesion as well as cell-cell and cell-medium friction, cell migration, cell division, and easily changeable elastic properties. Here, we have amended the original model by adding 1) the effect of a generic therapeutic radiation beam, 2) effective phagocytosis, and 3) the capability of simulating more than one cell type (here, healthy and cancer cells were used).
The effect of the radiation beam is implemented as a random process: it acts by picking a random number drawn for each cell. This number is then compared with the death probability . If the cells dies and cannot reproduce anymore. Other properties are left unchanged. This is essentially a simple Monte Carlo step.
II.2 Linear-Quadratic (LQ) model
The death probability () is taken from the Linear-Quadratic (LQ) model that was first introduced in the 1970’s [26, 27, 28]. The LQ model remains the most frequently used one [29, 30, 31, 32, 33, 8], and it is typically considered to work for doses in the range of 2-15 Gy. It assumes that the survival probability of a cell can be expressed as
where is the dose delivered by the beam, while () and (), called survival parameters, describe the radiosensitivity of the tissue. The larger these parameters, the more sensitive the cells are to radiation [34]. The parameter is considered to model a single-track process meaning that the cellular DNA is damaged in a single radiation event. The parameter describes a situation of a two-track process in which the DNA is damaged by two separate radiation hits. Thus, the ratio is an important characteristic of the model, providing the relative importance of single and double events.
The parameters and are experimentally accessible (albeit there can be important differences between in vitro measurements and in vivo outcomes) and they both depend on the nature of the beam and the target tissue. The rationale for Eq. 1 is that it is a mechanistic model describing radiation damage to the DNA caused by two different types of processes. More complex approximations for , including those based on the above LQ model, Lea-Catcheside factors, or extensions of the LQ model to low and high dosages are not considered here, but they can be easily implemented in the code. For generalizations of the LQ model, see, e.g., Refs. [31, 30, 33, 35, 36, 37].
II.3 The action of the radiation beam
The action of the beam during a time step is instantaneous. Once the program determines the set of dying cells, their status is switched from alive to dead during the same time step. In real treatments, a single dose delivery usually lasts from 10 to 30 minutes, and thus the instantaneous action may seem as a harsh approximation. This is, however, not the case due to time scales. In particular, the time unit in CellSim3D is determined by cellular growth [18]. The program saves the state of the system in the output file once every steps, where , namely the output frame frequency, is a number chosen by the user. Therefore, if the time interval between two successive frames reflects a period longer than 30 minutes, the beam can be considered as instantaneous.
In our simulations, the life cycle of a cancerous cell defined from the newborn state and ending with division, lasts less than 1,500 integration steps. In real time scales, the cell cycle of a human cancerous cell has a duration of about 24-72 hours, mitosis taking about 1 hour, setting one single frame to be in that range. Therefore, since one frame stands for a period longer than any dose delivery, cell death can be thought as instantaneous.
In reality, dead cells undergo phagocytosis [38]. We modeled this process as follows: when phagocytosis is activated, the program progressively lowers the internal pressure of the dead cells, until the threshold defined by the user is reached. The rationale is that by decreasing the volume of the dead cells they become very soft and flexible, and free up space for the live cells to proliferate. The effect of this is essentially the same as the full removal of the dead cells but without the computational complexity required by restoring the topology of the system and neighbor tables. Phagocytosis is, in general, much faster than cellular reproduction [38] and hence the shrinking rate for the dead cells is set to be ten times the growth rate of alive cells. The factor of ten is arbitrary but crucial (although its value can be changed if/when needed): while the unit of time for the simulation is determined by the cell cycle [25], it is important to keep the events that occur within a single simulation in the right order.
In the current approach, two different cell types are present simultaneously. The first cell type represents healthy and the second one cancerous cells; we note that the type of cells is not limited to two, but can be varied as necessary. The two types of cells are identical in all parameters but Young’s modulus, which describes the cell stiffness; cancer cells are typically considered to be softer than healthy cells [39]. Cells with a higher Young’s modulus are less prone to deformation and therefore grow slowly when compared to softer cells. Slower growth results in slower division, which mirrors the well-known fact that cancer cells proliferate faster than healthy ones [40]. In prior studies this approach has been successfully applied to study tumor growth in epithelial tissues [41].
III Results
System | for alive cells | for dead cells | for alive cells | for dead cells | % of alive cells |
Packed | - | - | 100 | ||
78.8 | |||||
60 | |||||
38.7 | |||||
Open | - | - | 100 | ||
80.3 | |||||
60.9 | |||||
39.5 |
The CellSim3D[18] model for mechanobiological studies of growth, division and proliferation of cells has been successfully validated against experimental data for the mitotic index and packing of epithelial cells [42], and it has been used to study tissue growth under different conditions [24, 43]. The addition of new features in CellSim3D requires two different types of validations. First, we compared the packing properties of cells to those obtained from experiments [44] for elastic spheres and found very good agreement. We compared also the contact forces and their distributions, in particular with respect to the jamming transition. The distribution of forces displays an exponential tail, in agreement with experimental data from elastic shells [44]. There are only very few force measurements in cellular systems, but Trepat et al. [45] measured traction distributions using Madin-Darby canine kidney cells, finding that the peak of the distribution can be well approximated by a Gaussian, while the tail is an exponential, in excellent agreement with our model predictions. This will be discussed in detail in Section III.1.1.
A second validation was performed with respect to radiation treatment, with two simulation setups. The first setup was aimed at testing the radiation beam in a common pathology such as prostate tumor: direct comparisons were made with published experimental results for radiation (photon) treatment for the human prostate cancer PC-3 cell line [46]. The second setup was aimed at the investigation of the force distribution in the system for a broad range of final states, varying the percentage of alive cells and the size of the box. We found good agreement despite the varying experimental setups. The main parameters of the simulations described in the next sections are summed up in Table 1. We will discuss these validations in the next two sections.
III.1 Contact force probability density
Eight different simulations were executed using different sizes of the simulation box, and percentage of dead cells at the final step. The spatial configurations are labeled as open in the case of an infinite box with open boundaries and packed in the case of a rigid box filled at the level of more than 80% at the final step of the simulation (confluence or close to confluence). The percentages of alive cells at the end of the simulation were, respectively, 100% (beam absent), 80%, 60% and 40%. We reserved, in each simulation, 6% of the time steps for the system to relax after radiation delivery. Up to 60,000 cells were simulated in each case.
It has been shown that near the jamming transition the probability density of the inter-particle forces decays nearly exponentially at large forces; the exponential tail is a signature of disorder and mechanical frustration. We observed that, as the packing density is increased, the tail of the distribution crosses over to a Gaussian indicating a more uniform force distribution. This is in line with previous experimental and simulation studies [44].
We extracted the probability density of the contact forces . The force is normalized by the mode of the distribution, such that the peak is at . The probability densities are renormalized after cutting off the last 1% of the distribution, since the points in that range have a very low statistics and represent, in the case of alive cells, newborns which have anomalously high values caused by the conformation after the division algorithm. We fitted the peak of the distribution with a Gaussian function
and the tail with an exponential function
in order to to determine the stress transfer and the state of the system with respect to the point of the jamming transition. In Eqs. 2 and 3 above, and are the respective amplitudes, the position of the peak and the width of the Gaussian.
III.1.1 Analysis of the fit
The parameter values from fitting to an exponential, Eq. 3, and the survival percentages are shown in Table 2, and the probability densities of the dead cells in the six simulations with the beam active are shown in Fig. 1. Table 2 shows interesting features: First, the steepness of the tails of the distributions is higher in the open systems than in the packed ones. This is expected, since the packed system is characterized by less space per particle, thus causing a more abundant population of stronger forces. The same steepness pattern is found when comparing dead and alive cells, in both spatial configurations. There are two interpretations:
The shrinking of the dead cells, i.e., mimicking phagocytosis, was not complete. Therefore, the majority of the dead cells have some free space around them and are not completely packed with their neighbors, resulting in softer contact and lower peak forces.
Implementation of cell shrinking: As a cell gains internal pressure when growing, it loses internal pressure when shrinking. Cells whose internal pressure is close to P, the lower threshold, have a weaker pressure force , which points from the center of mass outward, and, in principle, prevents the cell from folding. Cells with lower pressure are more prone to becoming concave in some areas, decreasing the contact area with the neighbors, and hence the total contact force on the cell.
We also observed that the relative errors of the dead cells are higher than those of alive cells. Moreover, the relative errors of the open space simulations are higher than those in the packed ones. These facts reflect the abundance of statistics at higher forces. In fact, the shorter the sample in the tail, the higher the probability for some bins of the histogram to contain only one or two points. The latter ones flatten the tail and augment the spread. This feature does not depend on the binning of the histogram and can be explained exactly as in the previous paragraph.
Parameter | Packed dead | Open dead |
The parameter values from fitting to a Gaussian, Eq. 2, are shown in Table 3. We notice that in the case of fixed volume (packed) spatial configuration, the peak is more than twice wider than in the case of free growth (open boundary) having and , respectively. This is easily explained by the fact that in a free growth environment only few cells at the very core feel strong forces from their neighbors, while cells at the border proliferate freely and tend to expand the system and reduce contacts with their neighbors rather than compress the system. This is not possible when the system has filled all the available volume, and higher forces are experienced throughout the system. This fact is also expressed in the statistical error of .
The exponential tails and the widths of the Gaussians provide insight on the state of the system with respect to the jamming transition. When the system has little to no volume available, cell mobility decreases dramatically, and the cells stick together forming a compact granular tissue. In this case, the force distribution has a long exponential tail with a very marked slope difference from the initial, Gaussian peak. This is evident in the force distribution of dead cells in the packed configuration in Fig. 1. A similar result can be found in Fig. 7b of [41], where the authors showed the force distribution of a system of bidimensional cells with high friction and no free volume available. On the other hand, when the system has room to expand at its borders and little free space in its core due to the deflation of dead cells, the exponential tail is still present but way steeper and almost contiguous to the Gaussian peak. This is the case of the dead cells in the open space configuration in Fig. 1. The system is still dense in its core, but cell mobility is still considerable and clusters of cells can move. This is again in agreement with [41], where the authors show in their Fig. 7a the same system but with null friction. Even if the volume available is still zero, the absence of friction facilitates independent cell movement and keeps the system at the jamming threshold.
III.1.2 Shape of the distributions
Figure 1 shows that the peaks of the probability distributions for the packed configuration are, in general, lower than the ones of the free distribution. This is simply caused by the longer tail, and the normalization of the distributions. Since the domain is longer and the area is fixed to unity, the peak at the mode must be lower. We notice that both tails have an exponential shape, but the cells that are in open space exhibit a steeper decay. Due to the normalization of the probability density, the cells in with the open boundary conditions are more peaked around the maximum. The difference in peak height is also found when comparing the force probability distributions of alive and dead cells in the packed vs. open graphs, as shown in Fig. 1. The shapes of the curves of cells from the same spatial configuration are basically the same.
Label | [Gy]-1 | [Gy]-2 | Data from |
-1 | 0.064 | 0.0167 | Deweese et al. [47] |
-2 | 0.241 | 0.067 | Algan et al. [48] |
-3 | 0.487 | 0.055 | Leith et al. [49] |
III.1.3 Comparison with experimental data
We now compare our results with the experimental data of Jose et al. [44]. They prepared thousands of micron-sized elastic shells suspended in a solvent, applied different loads to the samples, and extracted the force distribution. They found an initial peak and a long exponential tail. The most impressive similarities are found between our distributions in the packed configuration in Fig. 1, and their curves at low fractions of volume occupied by grain. This is expected, since the system in our simulation does not fill completely the volume at its disposal, but it is still compressed in five out of six sides of the cubic box.
There are, however, two notable differences. The first is that in our work, the gradient between the peak and the initial value, at , is high, ranging from two to three orders of magnitude. In their case, the difference is small and, in some curves, displays a plateau. Their abundance of low forces is explained by the fact that, in the small force domain, they compute the forces with a linear equation, starting from the deformation of the grain. They observe a plateau when the fraction of space occupied by the grains is lower than 0.7. The hard sphere packing theorem [50] states that the maximum volume occupied by hard spheres packed in space is , while the limit for random packing is [51].
Therefore, since the density of soft grains in Ref. [44] is even smaller than the limit density of hard grains, and it is comparable with the threshold of randomly packed hard spheres, multiple soft grains in their experiment must have no deformation at all, giving abundance of points at low force. On the other hand, in our case, cells proliferate from an initial core, and any cell at the border of the system has at least one neighbor with which it is in contact, thus yielding a force contribution close to the mean and far from zero.
The second difference to the distributions in Ref. [44] is the fact that they measured forces up to 4 times the mean force, while we detected values that were 7 times higher than our peak value (which is proximal to the mean and set to unity in Fig. 1). This fact, however, is explained by how the forces act In a packed configuration, neighbor cells are constantly pushed one against the other, and a linear repulsion arises. Since the force has no upper bound, the repulsion force can become arbitrarily large, and in close packing it can easily reach high values.
They also observed a Gaussian behavior away from the jamming point by augmenting the load, and therefore the fraction of space occupied by the grains. Although our open boundary simulations also show a behavior similar to a Gaussian right after the peak, the setup is too different to be reliable for a comparison, since there is no load on our system, and the packing in the core of the system is only due to cell-cell adhesion.
Trepat et al. [45] measured traction distributions in canine kidney cells. Although we cannot compare their results quantitatively – we have force distributions – qualitative features can be compared. The data and fits shown in Fig. 1 are in excellent agreement with the traction data of Trepat et al. [45] which also shows Gaussian peaks and exponential tails.
III.2 Cell survival: the comparison with prostate cancer data
Three data sets of survival parameters (, ) were extracted from the ratios present in literature [55] along with three data sets [52, 53, 54] on the dead-to-alive ratios for the PC-3 cell line. In all cases radiation was delivered by a photon beam with no fractionation. We list the survival parameters in Table 4.
The survival parameters and do not depend on the dose , but only on the nature of the beam and the structure of the tissue. However, these experimental values come from studies with different aims, namely the impact of the over-expression of a specific gene or a specific protein on the radiosensitivity of the tissue. Hence, these experimental data are influenced by external factors and the cell culture samples may also meet varying experimental conditions due to specific preparations, as listed below:
In Kiprianou et al. [52] the influence of an over-expression of a specific gene on radiosensitivity was carefully analyzed and compared with reference measurements performed on a pure PC-3 culture. Here, we used these unbiased reference data.
Russell et al. [53] investigated the radiosensitivity enhancement caused by the expression of a specific protein. The study presents a control, in which the sample was mixed with a specific concentration of dimethyl sulfoxide (DMSO) before irradiation. In this work, we used these unbiased data.
In the case of Yun et al. [54] the radiosensitivity enhancement was caused by the lack of a protein. Also in this case, we used the reference control, where the irradiated sample presents an endogenous level of the protein.
It is important to notice that, especially for improvements and accuracy, although we have chosen unbiased (or less biased) published data sets, differences in sample preparation, vector substance, and cell counting may complicate any quantitative comparison.
The simulation results and the experimental data extracted from references are plotted in Fig. 2. As expected, quantitative agreement is mixed: our results are mainly qualitative, since we collected parameters and values from previous studies having quite different experimental conditions. Quantitative agreement is restored at the tails of the distributions.
More specifically, we notice that the first set of parameters (blue curve in Fig. 2) slightly overestimates the experimental set taken from Ref. [52] (orange curve) by a maximum % at Gy, while the difference becomes lower than 8% in all other cases. The disagreement further decreases toward the tail, as expected, since more cells are killed by the increased dose. The second set (cyan) and the experimental data taken from Ref. [53] (magenta), on the other hand, display a very good agreement, meaning that the measurements from Ref. [53] are well described by the parameters fitted in [48]. This is impressive, since the radiation is from a different source (respectively, X- and -ray) and a vector substance was added to the sample in the experimental setup, potentially slightly altering its radiosensitivity. The last set (black) describes a very sensitive tissue and the simulation systematically underestimates the experimental data taken from Ref. [54] (green).
In each simulation the final number of cells was between and , thus providing sufficient statistics with standard deviation for all points. The experimental data were published without error bars, although the authors reported that all points represented mean values of at least three samples.
Figure 3 shows three spatial projections of the final two frames of the force distribution with open boundary configuration and with a Gy dose (see Table 1). The mass densities of alive cells are shown in blue and those of the dead cells in red. Clusters, or islands, of dead cells are clearly visible across the tumor after 2,000 steps of proliferation. The dead cells are pushed together by the division of the alive cells, which occurs even in the core of the system due to the free space releases by the dead cells.
IV Conclusions
We have proposed and analyzed a new tool for radiation oncology simulations by representing the tissue as a set of cells, rather than a continuum material, using the recently proposed CellSim3D mechanobiological cell model. We implemented the effects of radiation beam with a customizable dose, tissue sensitivity and fractionation as a stochastic Monte Carlo process. In a first set of runs, we measured the probability density of the inter-cellular forces, finding an excellent agreement with experiments in many aspects, and some minor differences due to the celluylar nature of the system. With these measurements, we were able to determine the state of the system with respect to the jamming transition, gives information on how far the cells can move inside the tumor. As a further proof of concept, we investigated the effect of radiation on prostate tumor model, finding good qualitative, but mixed quantitative agreement with experimental data. This is likely due to the varying experimental conditions present in literature. For further development, the software can be made to include spatial and temporal inhomogeneities for the dose and tissue radiosensitivity as well as time dependence for, e.g., the dose rate. In addition, a more elaborated version of the survival probability can be used in place of Eq. (1), and cell migration can be used to simulate further dynamic effects of cancers, such as metastasis. Such extensions would allow for more quantitative simulations of fractionated dose, inhomogeneous irradiation protocols, or even an inhomogeneous radiosensitivity of the tissue, and provide useful comparisons with experiments or even predictions for actual treatments.
Acknowledgements.
We thank Francesco Cordoni for a critical reading of the manuscript. MK thanks the Discovery and Canada Research Chairs Programs of the Natural Sciences and Engineering Research Council of Canada (NSERC) for financial support.References
- cdc [2019a] Leading causes of death in females, United States 2017 (2019a), accessed: April 5, 2020.
- cdc [2019b] Leading causes of death in males, United States 2017 (2019b), accessed: April 5, 2020.
- Blyuss et al. [2015] O. Blyuss, L. Koriashkina, E. Kiseleva, and R. Molchanov, Optimal placement of irradiation sources in the planning of radiotherapy: Mathematical models and methods of solving, Comput. Math. Methods Med. 2015, 1 (2015).
- Bryant et al. [2017] A. K. Bryant, M. P. Banegas, M. E. Martinez, L. K. Mell, and J. D. Murphy, Trends in radiation therapy among cancer survivors in the united states, 2000-2030, Cancer Epidemiol. Biomarkers Prev. 26, 963 (2017).
- Bernier et al. [2004] J. Bernier, E. J. Hall, and A. Giaccia, Radiation oncology: a century of achievements, Nat. Rev. Cancer 4, 737 (2004).
- Connell and Hellman [2009] P. P. Connell and S. Hellman, Advances in radiotherapy and implications for the next century: A historical perspective, Cancer Res. 69, 383 (2009).
- Streffer and Herrmannn [2012] C. Streffer and T. Herrmannn, A century of development in radiation biology. basic principles of targeted and efficient radiotherapy, Strahlenther. Onkol. 188 Suppl 3, 231 (2012).
- Jones and Dale [2019] B. Jones and R. G. Dale, The evolution of practical radiobiological modelling, Br. J. Radiol. 92, 20180097 (2019).
- Chen [2014] W.-Z. Chen, Impact of dose calculation algorithm on radiation therapy, World J. Radiol. 6, 874 (2014).
- Woon et al. [2016] Y. L. Woon, S. P. Heng, J. H. D. Wong, and N. M. Ung, Comparison of selected dose calculation algorithms in radiotherapy treatment planning for tissues with inhomogeneities, J. Phys. Conf. Ser. 694, 012024 (2016).
- Tambe et al. [2011] D. T. Tambe, C. C. Hardin, T. E. Angelini, K. Rajendran, C. Y. Park, X. Serra-Picamal, E. H. Zhou, M. H. Zaman, J. P. Butler, D. A. Weitz, J. J. Fredberg, and X. Trepat, Collective cell guidance by cooperative intercellular forces, Nat. Mat. 10, 469 (2011).
- Jalalimanesh et al. [2017] A. Jalalimanesh, H. S. Haghighi, A. Ahmadi, and M. Soltani, Simulation-based optimization of radiotherapy: Agent-based modeling and reinforcement learning, Math. Comput. Simul. 133, 235 (2017).
- Enderling and Rejniak [2013] H. Enderling and K. A. Rejniak, Simulating cancer: Computational models in oncology, Front. Oncol. 3, 233 (2013).
- Marcu and Harriss-Phillips [2012] L. G. Marcu and W. M. Harriss-Phillips, In silico modelling of treatment-induced tumour cell kill: developments and advances, Comput. Math. Methods Med. 2012, 960256 (2012).
- D’Andrea et al. [2016] M. D’Andrea, M. Benassi, and L. Strigari, Modeling radiotherapy induced normal tissue complications: An overview beyond phenomenological models, Comput. Math. Methods Med. 2016, 2796186 (2016).
- Bobadilla et al. [2017] A. V. P. Bobadilla, P. K. Maini, and H. Byrne, A stochastic model for tumour control probability that accounts for repair from sublethal damage, Math. Med. Biol. 35, 181 (2017).
- Liu et al. [2021] R. Liu, K. A. Higley, M. H. Swat, M. A. J. Chaplain, G. G. Powathil, and J. A. Glazier, Development of a coupled simulation toolkit for computational radiation biology based on Geant4 and CompuCell3D, Phys. Med. Biol. 66, 045026 (2021).
- Madhikar et al. [2018] P. Madhikar, J. Åström, J. Westerholm, and M. Karttunen, CellSim3D: GPU accelerated software for simulations of cellular growth and division in three dimensions, Comp. Phys. Comm. 232, 206 (2018).
- Cogno et al. [2024] N. Cogno, R. Bauer, and M. Durante, Mechanistic model of radiotherapy-induced lung fibrosis using coupled 3d agent-based and monte carlo simulations, Commun. Med. 4, 023129 (2024).
- Garcìa Garcìa et al. [2024] O. R. Garcìa Garcìa, R. Ortiz, E. Moreno-Barbosa, N. D-Kondo, B. Faddegon, and J. Ramos-Méndez, TOPAS-Tissue: A framework for the simulation of the biological response to ionizing radiation at the multi-cellular level, Int. J. Mol. Sci. 25, 10061 (2024).
- Jones [1997] H. W. Jones, Record of the first physician to see Henrietta Lacks at the Johns Hopkins Hospital: history of the beginning of the HeLa cell line, Am. J. Obstet. Gynecol. 176, s227 (1997).
- [22] Open source: https://github.com/softsimu/cellsim3d.
- Madhikar et al. [2021a] P. Madhikar, J. Åström, B. Baumeier, and M. Karttunen, Jamming and force distribution in growing epithelial tissue, Phys. Rev. Res. 3, 023129 (2021a).
- Mazarei et al. [2022] M. Mazarei, J. Åström, J. Westerholm, and M. Karttunen, In silico testing of the universality of epithelial tissue growth, Phys. Rev. E 106, L062402 (2022).
- Mkrtchyan et al. [2014] A. Mkrtchyan, J. A. Åström, and M. Karttunen, A new model for cell division and migration with spontaneous topology changes, Soft Matter 10, 4332 (2014).
- Kellerer and Rossi [1972] A. M. Kellerer and H. D. Rossi, Theory of dual radiation action, Curr. Top. Radiat. Res. Quart. 8, 85 (1972).
- Chadwick and Leenhouts [1973] K. H. Chadwick and H. P. Leenhouts, A molecular theory of cell survival, Phys. Med. Biol. 18, 78 (1973).
- Kellerer and Rossi [1978] A. M. Kellerer and H. H. Rossi, A generalized formulation of dual radiation action, Radiat. Res. 75, 471 (1978).
- Dale [1985] R. G. Dale, The application of the linear-quadratic dose-effect equation to fractionated and protracted radiotherapy, Br. J. Radiol. 58, 515 (1985).
- O’Rourke et al. [2008] S. F. C. O’Rourke, H. McAneney, and T. Hillen, Linear quadratic and tumour control probability modelling in external beam radiotherapy, J. Math. Biol. 58, 799 (2008).
- Brenner [2008] D. J. Brenner, The linear-quadratic model is an appropriate methodology for determining isoeffective doses at large doses per fraction, Sem. Radiat. Oncol. 18, 234 (2008).
- Santiago et al. [2016] A. Santiago, S. Barczyk, U. Jelen, R. Engenhart-Cabillic, and A. Wittig, Challenges in radiobiological modeling: can we decide between LQ and LQ-l models based on reviewed clinical NSCLC treatment outcome data?, Radiat. Oncol. 11, 67 (2016).
- van Leeuwen et al. [2018] C. M. van Leeuwen, A. L. Oei, J. Crezee, A. Bel, N. A. P. Franken, L. J. A. Stalpers, and H. P. Kok, The alfa and beta of tumours: a review of parameters of the linear-quadratic model, derived from clinical radiotherapy studies, Radiat. Oncol. 13, 96 (2018).
- McMahon [2018] S. J. McMahon, The linear quadratic model: usage, interpretation and challenges, Phys. Med. Biol. 64, 01TR01 (2018).
- Cordoni et al. [2021] F. Cordoni, M. Missiaggia, A. Attili, S. M. Welford, E. Scifoni, and C. La Tessa, Generalized stochastic microdosimetric model: The main formulation, Phys. Rev. E 103, 012412 (2021).
- Cordoni et al. [2022] F. G. Cordoni, M. Missiaggia, E. Scifoni, and C. La Tessa, Cell survival computation via the generalized stochastic microdosimetric model (GSM2); Part I: The theoretical framework, Radiat. Res. 197, 218 (2022).
- Missiaggia et al. [2024] M. Missiaggia, F. G. Cordoni, E. Scifoni, and C. L. Tessa, Cell survival computation via the generalized stochastic microdosimetric model (GSM2); Part II: Numerical results, Radiat. Res. 201, 104 (2024).
- Aderem and Underhill [1999] A. Aderem and D. M. Underhill, Mechanism of phagocytosis in macrophages, Annu. Rev. Immunol. 17, 593 (1999).
- Alibert et al. [2017] C. Alibert, B. Goud, and J.-B. Manneville, Are Cancer Cells Really Softer than Normal Cells?, Biol. Cell 109, 167 (2017).
- Hanahan and Weinberg [2011] D. Hanahan and R. A. Weinberg, Hallmarks of cancer: The next generation, Cell 144, 646 (2011).
- Madhikar et al. [2021b] P. Madhikar, J. Åström, B. Baumeier, and M. Karttunen, Jamming and force distribution in growing epithelial tissue, Phys. Rev. Res. 3, 023129 (2021b).
- Madhikar et al. [2020] P. Madhikar, J. Åström, J. Westerholm, B. Baumeier, and M. Karttunen, Coarse-grained modeling of cell division in 3d: influence of density, medium viscosity, and inter-membrane friction on cell growth and nearest neighbor distribution, Soft Mat. 18, 150 (2020).
- Mazarei et al. [2023] M. Mazarei, J. Åström, J. Westerholm, and M. Karttunen, Effect of substrate heterogeneity and topology on epithelial tissue growth dynamics, Phys. Rev. E 108, 054405 (2023).
- Jose et al. [2016] J. Jose, A. van Blaaderen, and A. Imhof, Random three-dimensional jammed packings of elastic shells acting as force sensors, Phys. Rev. E 93, 062901 (2016).
- Trepat et al. [2009] X. Trepat, M. R. Wasserman, T. E. Angelini, E. Millet, D. A. Weitz, J. P. Butler, and J. J. Fredberg, Physical forces during collective cell migration, Nat. Phys. 5, 426 (2009).
- Kaighn et al. [1979] M. E. Kaighn, K. S. Narayan, Y. Ohnuki, J. F. Lechner, and L. W. Jones, Establishment and characterization of a human prostatic carcinoma cell line (PC-3), Invest. Urol. 17, 16 (1979).
- Deweese et al. [1998] T. L. Deweese, J. M. Shipman, L. E. Dillehay, and W. G. Nelson, Sensitivity of human prostatic carcinoma cell lines to low dose rate radiation exposure, J. Urol. 159, 591 (1998).
- Algan et al. [1996] O. Algan, C. C. Stobbe, A. M. Helt, G. E. Hanks, and J. D. Chapman, Radiation inactivation of human prostate cancer cells: the role of apoptosis, Radiat. Res. 146, 267 (1996).
- Leith et al. [1993] J. T. Leith, L. Quaranto, G. Padfield, S. Michelson, and A. Hercbergs, Radiobiological studies of pc-3 and du-145 human prostate cancer cells: x-ray sensitivity in vitro and hypoxic fractions of xenografted tumors in vivo, Int. J. Radiat. Oncol. Biol. Phys. 25, 283 (1993).
- Cadart et al. [2014] C. Cadart, E. Zlotek-Zlotkiewicz, M. Le Berre, M. Piel, and H. K. Matthews, Exploring the function of cell shape and size during mitosis, Dev. Cell 29, 159 (2014).
- Scott and Kilgour [1969] G. D. Scott and D. M. Kilgour, The density of random close packing of spheres, J. Phys. D 2, 863 (1969).
- N. Kiprianou and Rhee [1997] D. B. N. Kiprianou, E. D. King and J. G. Rhee, bcl-2 over-expression delays radiation-indeuces apoptosis without affecting the clonogenic survival of human prostate cancer cells, Int. J. Cancer 70, 341–348 (1997).
- Russell et al. [2003] J. S. Russell, W. Burgan, K. A. Oswald, K. Camphausen, and P. J. Tofilon, Enhanced cell killing induced by the combination of radiation and the heat shock protein 90 inhibitor 17-allylamino-17- demethoxygeldanamycin: a multitarget approach to radiosensitization, Clin. Cancer Res. 9, 3749 (2003).
- Yun et al. [2003] E.-J. Yun, C.-J. Lin, A. Dang, E. Hernandez, J. Guo, W.-M. Chen, J. Allison, N. Kim, P. Kapur, J. Brugarolas, K. Wu, D. He, C.-H. Lai, H. Lin, D. Saha, S. T. Baek, B. P. Chen, and J.-T. Hsieh, Downregulation of human dab2ip gene expression in prostate cancer cells results in resistance to ionizing radiation, Cancer Res. 70, 2829–2839 (2003).
- Carlson et al. [2004] D. J. Carlson, R. D. Stewart, X. A. Li, K. Jennings, J. Z. Wang, and M. Guerrero, Comparison of in vitro and in vivo ratios for prostate cancer, Phys. Med. Biol. 49, 4477–4491 (2004).